OpenStudy (anonymous):

solve for x, log(3X^2-8)=2

OpenStudy (anonymous):

write in equivalent exponential form first \[\log(A)=B\iff 10^B=A\]

OpenStudy (anonymous):

oh its automatically 10

OpenStudy (anonymous):

if you don't see a base, yes if it is \(\ln(x)\) then the base is \(e\) if it is \(\log(x)\) then the base is 10

OpenStudy (anonymous):

is the answer 3.46?